please help me with this problem:Expres .15(with the bar at the top of 15)as a fraction in simplest form.
Answer by mmm4444bot(95)

Hello There:
In case you don't know, the bar indicates that the digits below it repeat forever.
Here is an algebraic method to find the fraction for 0.1515151515...
x = 0.151515...
Now, multiply x by 100 to get
100*x = 15.151515...
The reason we choose to multiply by 100 is so that the repeating parts will line up if we subtract the first equation from the second. In other words:
100*x = 15.151515...
x = 0.151515...
-----------------------------
99x = 15
Divide both sides by 99 to solve for x.
x = 15/99
The numerator and denominator are both divisible by 3, so we can reduce this fraction.
x = 5/33
~ Mark